The acceleration due to gravity on the surface is 20 ms-2.The time period of a simple Pendulum on the surface of planey.if the time period is 2 S on the surface of the earth will be :(Take g=ms-2)


1
Expert's answer
2021-09-14T09:35:40-0400

If the time period is 2 s on the surface of the earth, then the length of the pendulum is


"T_E=2\pi\sqrt{\frac {l}{g_E}}\space\rightarrow \space l=\frac{g_ET_E^2}{4\pi^2}."

The period of the pendulum on the planet will be


"T=2\pi\sqrt{\frac {l}{g}}=T_E\sqrt{\frac {g_E}{g}}=1.4\text{ s}."